# Merging Facts, Crafting Fallacies: Evaluating the Contradictory Nature of Aggregated Factual Claims in Long-Form Generations

### 08 Feb 2024 v1

@@ B.1 Human Evaluation of FActScore
Based on the definition of FActScore in Min et al. (2023a), we ask human annotators to check if each atomic fact can be supported by Wikipedia. Unlike the annotation of D-FActScore, we do not ask the annotators to consider entity ambiguity or the context where the atomic fact is extracted. We also do not restrict the annotators to verify the atomic fact using a subset of the Wikipedia of the linked entity. Instead, as long as an atomic fact can be verified with any Wikipedia page, the atomic fact should be considered Supported.
@@ Appendix C Automatic Evaluation
(2) Split the atomic facts into groups, where each group of atomic facts corresponds to one distinguishable biography. We also use text-davinci-003 in this step, where we provide 4-shot demonstrations to teach the LLM how to split the atomic facts into groups. Two of the four demonstrations contain multiple distinguishable biographies, and the atomic facts from each biography should be grouped together. The other two demonstrations do not contain multiple biographies, and all the atomic facts in the paragraph belong to the same group of atomic facts. We ask the LLM to insert "$=$" among atomic facts that belong to different groups of atomic facts. Two demonstrations are shown in Table 5.
@@
C.1 Entity Linking
We introduce some notations which will be used in the explanation. Let $\mathcal{A}_{y}$ the all the atomic facts, assume that the atomic facts are split into $N$ groups of atomic facts $\{\mathcal{A}_{y,1},\cdots,\mathcal{A}_{y,N}\}$. For each group of atomic facts $\mathcal{A}_{y,i}$, our goal is to find a Wikipedia entity $e_{k}$ with a corresponding Wikipedia page that the facts in $\mathcal{A}_{y,i}$ refer to. With a slight abuse of notation, we will use $e_{k}$ to refer to an entity and the corresponding Wikipedia page.
Last, D-FActScore can be computed by the percentage of the atomic facts that can be supported by the Wikipedia page of a corresponding entity, i.e.,
[…]
$*\displaystyle\quad\sum_{a\in\mathcal{A}_{y,i}}\mathbbm{1}[a\text{ is supportedby $e_{i}^{*}$}].$
Appendix D Citation Recall
D.1 Definition
Shortcoming of Citation Recall
D.2 Experiment Result
We show the results of citation recall in Table 6. We discuss previouisly that citation recall cannot handle the non-factual paragraphs due to entity ambiguity. However, this is only a problem if citation recall is high enough to falsely make the readers believe that the paragraph is well-supported by the retrieved documents, but is it non-factual?
Citation Recall
ChatGPT 56.65
Llama-chat-13b 57.77
Llama-chat-70b 72.63
Tulu-v2-13b-dpo 69.32
Vicuna-7b 55.44

### 23 Feb 2024 v2

@@ B.1 Human Evaluation of FActScore
Based on the definition of FActScore in Min et al. (2023a), each atomic fact is verified using the whole knowledge source $\mathcal{C}$. Following this definition, we ask human annotators to check if each atomic fact can be supported by Wikipedia. Unlike the annotation of D-FActScore, we do not ask the annotators to consider entity ambiguity or the context where the atomic fact is extracted. We also do not restrict the annotators to verify the atomic fact using a subset of the Wikipedia of the linked entity. Instead, as long as an atomic fact can be verified with any Wikipedia page, the atomic fact should be considered Supported.
Our paper suggests that FActScore, while not initially tailored for paragraphs featuring entity ambiguity, may not fully capture the nuances in the evaluations of paragraphs outlined in Section 2. Our contribution seeks to identify and address this area for improvement within the original framework of FActScore, as described in Min et al. (2023a), with the goal of refining the metric to more adeptly handle diverse content scenarios.
B.2 Human Evaluation of Number of Distinct Entities
To evaluate the number of distinct entities by human evaluation, we give the human annotators the atomic facts decomposed from the paragraph. For each fact, we ask the annotators to use the Wikipedia pages we provide, which consists of the Wikipedia pages of the entities with the target name, to determine which entity the given atomic fact refers to. In this step, the annotators are instructed not to consider the context of the atomic fact, but consider the atomic fact as a standalone fact and find which entity this standalone atomic fact refers to. After we obtain the entity of each atomic fact in the paragraph, we can count the number of unique entities for the atomic facts in the paragraph; this number is the number of distinct entities.
@@ Appendix C Automatic Evaluation
(2) Split the atomic facts into groups, where each group of atomic facts corresponds to one distinguishable biography. We also use text-davinci-003 in this step, where we provide 4-shot demonstrations to teach the LLM how to split the atomic facts into groups. Two of the four demonstrations contain multiple distinguishable biographies, and the atomic facts from each biography should be grouped together. The other two demonstrations do not contain multiple biographies, and all the atomic facts in the paragraph belong to the same group. We ask the LLM to insert "$=$" among atomic facts that belong to different groups of atomic facts. Two demonstrations are shown in Table 7.
@@
C.1 Entity Linking
C.2 Using ChatGPT to Extract and Group Atomic Facts
The automatic evaluation of D-FActScore relies on using GPT3.5 to extract and group atomic facts. However, GPT3.5 has recently been deprecated (in mid Jan. 2024). Here, we show that all our experiment results hold when replacing GPT3.5 with ChatGPT; i.e., we use ChatGPT to extract atomic facts from the paragraph and use ChatGPT to split atomic facts into groups. We show the results in Table 5. Comparing the result of using ChatGPT (Table 4(c)) and using GPT3.5 (Table 4(b)), we find that while the absolute numbers slightly changes, the observations stated in Section 5.2 still holds. We recap those results here:
  • Using ChatGPT or GPT3.5 in automatic evaluation shows the same factuality rankings among ChatGPT, Llama-13b-chat, and Tulu-v2-dpo as the ranking obtained with human annotation.
  • D-FActScores obtained using ChatGPT and GPT3.5 are higher than human evaluation.
  • Automatic evaluation can determine the number of biographies accurately.
  • Open-source models lag behind ChatGPT.
  • ChatGPT fully uses the received documents.
Model FS D-FS # bio # ent.
ChatGPT 98.3 92.1 2.2 2.3
chat-13b 94.8 78.4 1.0 1.7
Tulu 91.9 83.2 1.3 1.7
(a) Human evaluation
Model FS D-FS # bio # ent.
ChatGPT 98.7 96.3 2.2 2.3
chat-13b 95.3 86.4 1.1 1.5
Tulu 95.8 88.5 1.3 1.7
(b) Automatic evaluation (GPT3.5)
Model FS D-FS # bio # ent.
ChatGPT 98.2 92.8 2.0 2.3
chat-13b 96.0 87.1 1.0 1.5
Tulu 96.3 88.6 1.2 1.7
(c) Automatic evaluation (ChatGPT)
Table 5: FS: FActScore, D-FS: D-FActScore, # bio: average number of separable biographies in one paragraph, # ent.: average number of distinct entities in one paragraph, chat-13b: Llama-13b-chat. Human and automatic evaluations are conducted on the same set of paragraphs. The result in Table 4(b) is obtained by using GPT3.5 to extract the atomic facts and split atomic facts into groups, and Table 4(c) is the result of using ChatGPT to extract the atomic facts and split the atomic facts into groups.
Appendix D Citation Recall
D.1 Definition
Shortcoming of Citation Recall
D.2 Experiment Result
We show the results of citation recall in Table 6. We discuss previouisly that citation recall cannot handle the non-factual paragraphs due to entity ambiguity. As a result, even if the citation recall is high, the paragraphs can still be non-factual. However, we do not see such a problem in the paragraphs generated in Section 2. This is because the citation recalls of the paragraphs generated by the LLMs are not very high, so there is no such a problem like "high citation recall but non-factual." Nevertheless, we stress that citation recall cannot handle non-factual paragraphs due to entity ambiguity.
Citation Recall
ChatGPT 56.65
Llama-chat-13b 57.77
Llama-chat-70b 72.63
Tulu-v2-13b-dpo 69.32
Vicuna-7b 55.44